Four capacitors are connected in a circuit as illustrated. Find the charge on, the voltage difference across, and the energy stored for each of the capacitors.

tor? Four capacitors are connected in a circuit as 57. .. t as illas trated in ▼ Fig. 16.29. Find the charge on, thevas inu the ference across, and the energy stored for each of capacitors C1 0.40 μF 0.40 μF C3 CA 12 V 0.20 μF 0.60 μF
Answer #1

Capacitors C1 and C2 are in parallel. Hence their net capacitance C12 = (0.40 + 0.40) \muF = 0.80 \muF

Similarly, Net capacitance of C3 and C4 in parallel, C34 = (0.20 + 0.60) \muF = 0.80 \muF

Thus we have two equal capacitances C12 and C34 connected in series across 12 V supply. Hence potential difference acros each capacitance C12 and C34 is 6 V. As C1 and C2 are in parallei and C3 and C4 in parallel, therefore

P.D. across C1 = P.D. across C2 = P.D.across C3 = P.D. across C4 = 6 V

Charge on C1 = Charge on C2 = (0.4 \muF) x 6 V = 2.4 \muC

Charge on C3 = (0.20 \muF) x 6 V = 1.2 \muC

Charge on C4 = (0.60 \muF) x 6 V = 3.6 \muC

Energy stored in C1 = Energy stored in C2 = (CV2) / 2 = (0.40 \muF) x 36 V 2 / 2 = 7.2 \muJ

Energy stored in C3 = (CV2) / 2 = (0.20 \muF) x 36 V 2 / 2 = 3.6 \muJ

Energy stored in C4= (CV2) / 2 = (0.60 \muF) x 36 V 2 / 2 = 10.8 \muJ